What was Real GDP in 2008 (in 2009 $)? ___________ Question 29 options: $11.48 trillion $14.25...

Formula 4

What was Real GDP in 2008 (in 2009 $)? ___________

Question 29 options:

$11.48 trillion

$14.25 trillion

$14.35 trillion

$17.81 trillion
Answer #1

The answer is "14.25 Trillion $"

2008

Nominal GDP = 14.3 Trillion $

CPI = 215.3

Real GDP = NOmianl GDP * 100 / CPI

= 14.3 * 100 / 215.3 = 6.6419 Trillion $

2009

CPI = 214.5

Real GDP in 2009 dollars = Real GDP in 2008 * CPI in 2009 / 100

= 6.6419 * 214.5 / 100

= 14.25 Trillion $